Christophe, Cyril, Daniel, Denis, Julien, Philippe, and Pierre (+ Fred in Calern) & the CHARA team & the rest of the SPICA team

N3 UT2022-03-12

  • Good conditions tonight (r0 7-8 cm). Thanks to Narsi and Norm, we tuned a little bit more the AO on S1 and S2 and the difference in the correction in the visible has been very clear. The minimum is clearly to use the fastest rate, the CL link, to tune the gain, and to optimize the alignment of the WFS. At least these steps are mandatory before considering other improvements.
  • We did some tests of fiber explorer + tip:tilt correction but we need to upgrade our control software to be more automatic and precise on this critical step of the observation.
  • We encountered a strange behaviour of the science detector, that seems to lost its sensitivity or power of intensification. Its cooling was also strange and at the end the images were really poor on the camera. It has been decided to remove the camera and to bring it back to France for maintenance.

D19 11 Mar 2022

  • Cleaning of lab, adjustment of air circulation for the science detector
  • Software activities for installing (difficultly) the SPICA station in the control room.
  • New tests of the cooling of the science detector: without the chiller the temperature can not go below -66°C which is quite high (nominal temperature should be -90°C). The control detector is stabilized (without the chiller at -75°C). This difference is probably due to the smallest circulation of air around the science detector. Adjustements may be needed in the future, depending on the performance with the additional chiller cooling.

N2 UT2022-03-11

  • High winds and very poor seeing prevented us to do anything useful on sky. Software debug on fast tip/tilt corrections and real-time display were done, partly on star and mostly on STS.

D18 10 Mar 2022

  • Installation of the hot-air evacuation of the two detectors
  • Final installation of SPICA software on the lab laptop. Preparation of the control room station.
  • Replacement of the FBI translation stage by a one using a stepper motor for standarization purpose.

N1 UT2022-03-10

  • humidity kept us closed for the first hours of the night. At UT7h00 we slewed to bright stars and get all scopes and AO aligned. 3 beams were perfectly aligned in SPICA but three were not present due to an error in the Coudé alignment of CHARA. After correction we got the 6 stars and the 6 pupils. r0 was arounf 6-7cm, seeing of the order of 1.5'', very fast. We got flux in the fibers but with a very low coupling estimated at the order of less than 1% of the light. On this bright star, MIRCx fringes on S1S2 were barely present (SNR of the order of 1) despite an Hmag of 2.3. Some blind recording were done on the control and on the science detector. We clesed early as humidity was again around 80% and seeing very fast.

D17 9 Mar 2022

  • Early morning test with Fred of the new server for the science detector and the associated RTD client. This is working very well now despite some difficulties related to the difference of axis between MIRCx and SPICA (spatial and spectral).
  • Main activities today around understanding our problems with PDC and the defaults of the visible dichroics. Even if we finally found the correct position of PDC2 for fringes B1B2, it turns that on fringes B1B3 the birefringence of the dichroics is really perturbating the measurement of contrast. At the end we decided to stop optimizing the PDC because we are starting to compensate the birefringence of the dichroics with the PDC. And for the first on-sky tests, we decided to add a polarizer at the output of the V-groove.
  • We measured the flux in visible and IR with a ND1 replacing one of the visible dichroic. Calculations to be made...
  • Finally, thanks to the real-time display provided by Fred (thanks a lot), we optimized the contrast of the smallest fringes by a huge factor by rotating the V-groove along its axis. All fringes, even the B1B6 ones, are now looking fine. We reoptimized the centering of the interferometric channel on the detector and redo the kappa matrix.

D16 8 Mar 2022

  • Installation of the laptop with the SPICA clients. Configuration (by Chris and Nils) of the control room machine for the SPICA observations. 
  • Correction of the motion of the PDC for a better reliability of their positioning
  • We started to evaluate the possibility of moving the PDC with the control of a polarizer and checking the opd variations fro mone polarization to another but the method is not sensitive enough to reliably evaluate the positions. So we moved to data recording and data analysis.
  • First tests of PDC calibration by using PDC1 at 70° and moving PDC2 by steps to move in the direction of the best contrast.pngB1B2 (not yet finished).
  • We tested the optics of the STS to identify why some of the beams exhibit absorption lines.The first test was done by changing one of the VIS-FOLD dichroic by a mirror and the problem disappeared. It is thus clear that these ots dichroics are of very poor quality. With the issue also encountered on the very low transmission of the green laser in the IR it is thus clear that we will have to change them.

D15 7 Mar 2022

  • Activities on the spectral calibration in MR, LR. 
  • First recording of kappa matrix
  • New adjustments of the position of beams on the detector after the dispersion and fine tuning of the direction of the dispersion in the different resolutions
  • End of day on kappa matrix again (after the fine adjustments) and we started the PDC calibration. 

D14 6 Mar 2022

  • Arrival of Philippe and Cyril. 
  • Continuation of the work on the alignment templates and data recording on detectors

D13 5 Mar 2022 SPICA is alive! pdfDay_13.pdf

  • Alignment of the spectrograph
    • Alignment of the mirror at 0 order
    • Adjustment of the cylindrical lenses
    • Alignment of the photometric channels on a vertical column
    • Alignement of the orientation of the dispersion in the different resolution
    • And finally fringes in all spectral resolution
  • Consolidation of software and electronics

D12: 4 Mar 2022 

  • Validation of the control software of the Differential Delay Lines, consolidation of the management of the servers and of the multivolt communication.
  • Finalization of the covers of the spectrograph table.
  • Alignment templates
    • we started with the reference template taking the position of reference of the FBI in image and pupil plane.
    • we used these references to do the automatic alignment of the beams coming from the STS after the installation of the ADC and PDC. 
    • The interaction matrix template has some issues, corrected remotely by Frédéric. 
    • TTT matrix done also.
    • More work needed to totally validating the alignement templates.
  • Start of the alignment of the spectrograph
    • Without the cylindrical lenses, orientation of the V-groove to go at the right position on the mirror position. First adjusment of the photometric channel to correctly match, together with the interferometric channel the entrance of the camera lens.
    • Orientation of the dispersion mechanism to have the non-dispersive position correctly oriented, and centered on the column 512 of the detector.
    • Addition of a F700 lens to reimage the output of the 6 fibers on the detector. Correction of the orientation of the V-groove to have the 6 fibers perfectly aligned on the columns of the detector
  • Departure of Christophe and some good time in Glendale downtown in the afternoon.jpgSafe travel Christophe

D11: 3 Mar 2022 pdfDay_11.pdf 

  • ADC have been installed on the periscope with the associated cabling. 
  • PDC are now operational and included in the SPICA software (movsmall movie)
  • Today we finalized the co-alignment of SPICA and CHARA
    • Improvement of the alignment of STS and CHARA beams
    • FBI beams adjusted on the STS beams on the bottom mirrors of the periscope by adjusting the M2 mirrors.
    • STS beams adjusted on the FBI beams at the entrance of SPICA table by adjusting the M1 mirrors of the periscope
    • Control of the image and pupil planes of the STS beams on the control detector and fine adjusment of the images position.
    • Then installation of the field lenses and adjustement of their XYZ position to get the STS pupils at the position of the FBI pupils and the size of the STS pupils equal to their theoretical value of 54 pixels.
    • The next step now will be the spectrograph.
  • Installation of the pipes for the water cooling of the two andor detectors
  • Continuation of the cleaning of the cabling
  • Theo showed us how to get the air thickness in OPLE: this will be useful for the initial manual settings of the LDC before the installation of their automatic positioning.

D10: 2 Mar 2022 pdfDay_10.pdf

  • Successfull morning session with Sylvain for the tests of the new version of SPICA-FT with the STS.
  • Installation of variable LEDs below the covers of the injection and spectrograph tables. 
  • The multivolt box exhibits always some issues; It was finally found that the current is probably to high for the 5 controllers of the delay lines. We decided to dispatch the 5 delay lines on 3 switches to balance the current. It works! To be checked more in the coming days.
  • The software control of the ADC is finally fixed. Both prisms rotate now in the same direction, as decided in the convention, and they are able to move on the shortest path. Angles are defined from -180° to +180°, 0° directed towards Zenith and positive towards East. The intialisation put them all at 0° corresponding to no dispersion compensation.
  • Fine alignment and focusing of the control beams.
  • Adjustment of the 6 images of the FBI on the control detector by tiny motions on the beamsplitters.

D09: 1 Mar 2022 pdfDay_9.pdf

  • Morning session with Sylvain in Nice for SPICA-FT tests. Progresses on the communication with OPLE but no way this morning to use MIRCX+STS because of issues on the Credone, solved in the afternoon by Narsi.
  • A small increase of the voltage of the Multivolt box has finally solved the issue.
  • Continuation of the alignment of the table
    • Check of the centering of the beam on the exit lens of the injection module.
    • Orientation of the injection modules to center the beams on the Fast TT mirrors at 5V, and control of the centering on the collimator tube. 
    • Adjustment of the support of the FTT to align the beams on the PDC support without the focuser.
    • Insertion of the focuser and alignment of the beams on the periscope masks.
    • Adjust the focus of the focuser and realignment on the periscope.
  • Alignment of the control beams
    • orientation of the beam splitters to center on M4
    • Global adjustment of M4s to go on the parabola.
    • Adjustement of the centering of the corner cubes on the beams. Iteration with CC and BS.
    • Adjustment of the parabola to send the beams on the M5 mirror in front of the camera.
  • Continuation of the work on the ADC motions. Almost done hopefully but we were not able to finally check because of the start of the night.
  • Validation of the CHARA low-level software to drive the Visible LDC. 
  • The SPICA computer in the control room with 4 screens is almost ready. 
  • After the work in lab, we work on the preparation of the LEDs that will permit lightening the tables.

D08: 28 Feb 2022 pdfDay_8.pdf

  • Installation of the limit switchs on the STS VIS-FOLD for the security of the motor.
  • Continuation of the software for ADC; cabling of the power of the shutters. Functional tests of the dispersion mechanism.
  • Installation of the additional fibers from the V-groove to the top of the cover of the injection table (alignment sources, spectral lamps).
  • Alignment of the FBI
  • Installation of the 6 main fibers from the V-groove to the injection devices
    • alignment of the polarization axis of the 6 fibers with 2-crossed polarizing plates, using the rotation along their main axis of the injection tubes
    • focalization of the 6 injection modules with the 730nm laser and a shear plate.
  • Installation of the isolated pipes between the location of the chiller in the cabinet and the detectors on the tables.
  • Diagnostic of the problems on the multivolt box.

D07: 27 Feb 2022 pdfDay_7.pdf

  • Continuation of the preparation of the fiber protection between the two tables. Final positioning of the spectrograph table.
  • Continuation of the cabling and testing of all components. Everything is in order, we just need to finalize the testing of the dispersion mechanism. 
  • Continuation of the software control for the ADC.
  • Final cabling and testing of the STS, thanks to the help of Narsi. Its control is now integrated in the shutter GUI, together with the IR FOLD. Preparation of the addition of two limit switches for a better protection of the motor.
  • Preparatory work on SPICA-FT in parallel.
  • The installation is now ready to start the optical alignment of SPICA-VIS.

D06: 26 Feb 2022 pdfDay_6.pdf

  • Continuation of the work on the spectrograph table, and cabling of all the components between the two tables. Installation of the FBI (Fiber Back Illumination), and adaptation of the cover for the protection of the fibers.
  • Finalization of the power server & power gui but one issue with the electronics of the Multibox (multiple switch). Installation of the Andor Control Camera with its hot-air evacuation system.
  • Continuation of the tests of the ADC
  • New cabling for the STS VIS FOLD and test of software with Narsi
  • End of afternoon in Pasadena

D05: 25 Feb 2022 pdfDay_5.pdf

  • Installation of the spectrograph table at its location and preliminary adaptation of the cover.
  • Installation of the spectrograph modules
  • Continuation of the installation of the electronics and cables
  • Testing the power server and the power gui: almost everything is working fine (with power in 110V of course...)
  • The ADC motors are now working correctly, after much difficulties encountered in Nice during the (very...) long period of tests!

D04: 24 Feb 2022 pdfDay_4.pdf

  • Installation of almost all the modules (but the PDC and the pupil lens) on the table. Cabling, and continuation of the installation of the electronics and all the cables (injection table, periscope table)
  • Tests of the IMG mirrors (the one on top of the periscope) 
  • Test of almost all the remotely controlled power plugs.  
  • Tests of various pieces of software for the control of modules.
  • Installation of the ADC devices for finalization of their control software.
  • Short afternoon as CHARA is starting again the Array for preparing the new observing season.

D03: 23 Feb 2022  pdfDay_3.pdf  (additional pictures from Julien jpgIMG_1014.jpg jpgIMG_1015.jpg jpgIMG_1043.jpg jpgIMG_1046.jpg jpgIMG_1047.jpg jpgIMG_1060.jpg)

  • Final alignment of the SPICA table with respect to the CHARA beams
    • installation of the VLDC and control of the potential deviation of the reference beam on the far target. Nothing critical to note, which confirms the correct optical assembly of the LDC.
    • removing the supports of the VIS-FOLD to lower the dichroïcs and reduce the risks of vignetting the IR beams. Realignment of the VIS-FOLD from the target on the metrology table to the target at the entrance of the periscope. Finally  we conclude that some misalignments of the STS are not possible to compensate. At the end we have some small and non really critical defaults on the STS pupils in the Visible beams.
    • Fine adjustment of the mirrors of the periscope on the STS beam to correctly match the targets on the CHARA visible table, on beams 1 & 6. The targets on the SPICA table are very close to the beams but not perfect. At the end we concluded that the CHARA visible table and the SPICA table are not coplanar. We finally decided to more the upper mirrors of the periscope to correctly match the SPICA targets. A small correction of the heights of the North feet of the SPICA table permitted to correctly match the close and far targets at the same time and on beams 1 and 6. This step closes the positioning of the injection table.
  • Installation of all the cables from the rack to the tables, and installation of the electronics module on the cover. The two ANDOR detectors have been started for remote activities.
  • Start of the installation of the modules on the table. But it appears that the alignment of the holes on the table is very poor with sometimes almost 0.5mm of error. Therefore the installation with the pins and the screws is much more complicated than foreseen. Some holes for the fixation screws will need to be slightly enlarged.
  • All devices are now correctly set on their IP address and all the codes have been adapted.

D02: 22 Feb 2022 (pdfDay_2.pdf, movmovie of the periscope)

  • Fine positioning of the table with respect to the CHARA visible table. Adjustement of the cover and adaptation of the evacuation of the hot air of the Control Detector.
  • The periscope is now operational and it has been installed at its location.
  • Preparation of the motion of the VIS-FOLD for the alignment. 
  • Configuration of all the electronic devices with their final IP, their final cables.
  • Cabling of the rack and functional tests.The three main servers are operational.
  • Validation of the coding of the remote-power-plug server, at least for all the active devices of the rack.
  • Organisation of the toolbox and of the SPICA shelf.

D01: 21 Feb 2022 (pdfDay1_pdf)

  • Mechanical disassembling of the periscope; installation of the Thorlabs mirrors
  • Gluing of the last LDC prism, alignment of the last LDC, pre-installation at its final location
  • Mounting and installation of the cover of the injection table
  • Installation of all the components in the SPICA cabinet
  • Final cabling of the electronic boxes